# Andreas Schlaginhaufen SYCAMORE, EPFL

### v1

@@ 1 Introduction
  • We experimentally validate our results in a gridworld environment (Section 5).
@@ 2 Background
Given $x$ and $y$ in some Euclidean vector space $\mathcal{V}$, we denote the $p$-norm by $\|x\|_{p}$, the orthogonal projection onto a closed convex set $\mathcal{X}\subset\mathcal{V}$ by $\Pi_{\mathcal{X}}(x)=\arg\min_{y\in\mathcal{X}}\|x-y\|_{2}$, and the standard inner product by $\langle x,y\rangle$. For a linear operator $A$, we denote its image and rank by $\operatorname{im}A$ and $\operatorname{\operatorname{rank}}A$, respectively. Given two sets $\mathcal{X}$ and $\mathcal{Y}$, we denote $\mathcal{X}+\mathcal{Y}$ for their Minkowski sum and $\mathcal{Y}^{\mathcal{X}}$ for the set of all functions mapping from $\mathcal{X}$ to $\mathcal{Y}$. Additionally, we denote $\Delta_{\mathcal{X}}$ for the probability simplex over $\mathcal{X}$ and $\mathbbm{1}$ for the indicator function. The interior $\operatorname{\operatorname{int}}\mathcal{X}$, the relative interior $\operatorname{\operatorname{relint}}\mathcal{X}$, the relative boundary $\operatorname{\operatorname{relbd}}\mathcal{X}$, and the convex hull $\operatorname{\operatorname{conv}}\mathcal{X}$ of some set $\mathcal{X}$ are defined in Appendix A, along with an overview of all other notations.
@@
We consider a regularized MDP (Geist et al., 2019) defined by a tuple $\left({\mathcal{S},\mathcal{A},P,\nu_{0},r,\gamma,h}\right)$. Here, $\mathcal{S}$ and $\mathcal{A}$ represent finite state and action spaces with $|\mathcal{S}|,|\mathcal{A}|>1$, $\nu_{0}\in\Delta_{\mathcal{S}}$ the initial state distribution, $P\in\Delta_{\mathcal{S}}^{\mathcal{S}\times\mathcal{A}}$ the transition law, $r\in\mathbb{R}^{\mathcal{S}\times\mathcal{A}}$ the reward, and $\gamma\in(0,1)$ the discount factor. Furthermore, $h:\mathcal{X}\to\mathbb{R}$ is a strictly convex regularizer that is defined on a closed convex set $\mathcal{X}\subseteq\mathbb{R}^{\mathcal{A}}$ with $\operatorname{\operatorname{relint}}\Delta_{\mathcal{A}}\subseteq\operatorname{\operatorname{int}}\mathcal{X}$. The goal is to find a Markov policy $\pi\in\Delta_{\mathcal{A}}^{\mathcal{S}}$ maximizing the regularized objective $\mathbb{E}_{\pi}\left[{\sum_{t=0}^{\infty}\gamma^{t}\left[{r(s_{t},a_{t})-h\left({\pi(\cdot|s_{t})}\right)}\right]}\right]$. Following the classical linear programming approach to MDPs (Puterman, 2014), this can be cast equivalently as the convex optimization problem
@@
Due to the strict convexity of $h$, the regularized MDP problem has a unique optimal policy (Geist et al., 2019), hence guaranteeing the uniqueness of the optimal occupancy measure in (O-RL). In addition, we assume that the gradients of $h$ become unbounded towards the relative boundary of the simplex as detailed in Assumption 2 below. {restatable}[Steep regularization]assumptionsteepregularization Suppose that $h:\mathcal{X}\to\mathbb{R}$ is differentiable in $\operatorname{\operatorname{int}}\mathcal{X}$ and that $\lim_{l\to\infty}\left\lVert\nabla h(p_{l})\right\rVert_{2}=\infty$ if $\left({p_{l}}\right)_{l\in\mathbb{N}}$ is a sequence in $\operatorname{\operatorname{int}}\mathcal{X}$ converging to a point $p\in\operatorname{\operatorname{relbd}}\Delta_{\mathcal{A}}$. Assumption 2 ensures that the optimal policy is non-vanishing, and together with Assumption 2 below, we also have that the optimal occupancy measure is non-vanishing. {restatable}[Exploration]assumptionocclowerbound Let $\nu(s)\vcentcolon=\sum_{a}\mu(s,a)\geq\nu_{\text{min}}>0$ for any $s\in\mathcal{S}$ and $\mu\in\mathcal{M}$. One way to guarantee Assumption 2 is to impose a lower bound on the initial state distribution $\nu_{0}$. In the following, it will be convenient to denote the optimal solution to (O-RL) for the reward $r$ as
@@
$\ell(r,\mu)\vcentcolon=\max_{\mu^{\prime}\in\mathcal{M}}J(r,\mu^{\prime})-J(r,\mu).$ (3)
That is, $\mu=\mathsf{RL}(r)$ if and only if $\ell(r,\mu)=0$.
@@
As we aim to analyze the transferability of rewards to new transition laws $P\in\Delta_{\mathcal{S}}^{\mathcal{S}\times\mathcal{A}}$, it will often be useful to explicitly specify the dependency on $P$. We do so by adding a subscript – e.g. we write $\mathcal{M}_{P}$, $\mathsf{RL}_{P}$, and $\ell_{P}$. However, for better readability, we drop these subscripts whenever there is no potential for confusion.
@@
However, since we don’t have direct access to the expert’s policy but only to a finite set of demonstrations, the best we can hope for is an algorithm that with high probability outputs a reward $\hat{r}\in\mathcal{R}$ such that $\ell(\hat{r},\mu^{\textsf{E}})$ is small – i.e. an algorithm that is PAC (Syed and Schapire, 2007).
@@
The reward corresponding to an optimal occupancy is not unique. For example, all rewards in the affine subspace $r+\mathcal{U}$, where $\mathcal{U}\vcentcolon=\operatorname{im}(E-\gamma P)$ is the subspace of so-called potential shaping transformations, correspond to the same optimal occupancy measure (Ng et al., 1999). Hence, it is often convenient to consider these rewards as equivalent (Kim et al., 2021) and to measure distances between rewards in the resulting quotient space. Given a linear subspace $\mathcal{V}\subset\mathbb{R}^{\mathcal{S}\times\mathcal{A}}$, the quotient space $\mathbb{R}^{\mathcal{S}\times\mathcal{A}}/\mathcal{V}$ is the set of all equivalence classes $[r]_{\mathcal{V}}\vcentcolon=\left\{{r^{\prime}\in\mathbb{R}^{\mathcal{S}\times\mathcal{A}}:r^{\prime}-r\in\mathcal{V}}\right\}$. We endow $\mathbb{R}^{\mathcal{S}\times\mathcal{A}}/\mathcal{V}$ with the quotient norm $\left\lVert[r]_{\mathcal{V}}\right\rVert_{2}\vcentcolon=\min_{v\in\mathcal{V}}\left\lVert r+v\right\rVert_{2}=\left\lVert\Pi_{\mathcal{V}^{\perp}}r\right\rVert_{2}$ and we say that $r$ and $r^{\prime}$ are close in $\mathbb{R}^{\mathcal{S}\times\mathcal{A}}/\mathcal{V}$ if $\left\lVert[r]_{\mathcal{V}}-[r^{\prime}]_{\mathcal{V}}\right\rVert_{2}$ is small. Moreover, the expert’s reward is said to be identifiable up to some equivalence class $[\cdot]_{\mathcal{V}}$ if $\mathsf{IRL}(\mu^{\textsf{E}})\subseteq[r^{\textsf{E}}]_{\mathcal{V}}$. In this paper, we will consider the equivalence relations induced by constant shifts, i.e., $\mathcal{V}=\bm{1}\vcentcolon=\left\{{r\in\mathbb{R}^{\mathcal{S}\times\mathcal{A}}:r(s,a)=c\in\mathbb{R}}\right\}$, and by the aforementioned potential shaping transformations, i.e., $\mathcal{V}=\mathcal{U}$. Note that since $\bm{1}$ is a subspace of $\mathcal{U}$ and $\mathcal{U}$ is $|\mathcal{S}|$-dimensional, $[r]_{\bm{1}}$ is a strict subset of $[r]_{\mathcal{U}}$ whenever $|\mathcal{S}|>1$.
@@ 3.1 Problem formulation
consisting of trajectories sampled independently from the experts $\mu^{\textsf{E}}_{P^{0}},\ldots,\mu^{\textsf{E}}_{P^{K-1}}$. Each expert is optimal for the same unrevealed reward $r^{\textsf{E}}\in\mathcal{R}$, but under different transition laws, $P^{0},\ldots,P^{K-1}$. Our goal is to recover a reward $\hat{r}\in\mathcal{R}$ that is transferable across a set of transition laws $\mathcal{P}\subseteq\Delta_{\mathcal{S}}^{\mathcal{S}\times\mathcal{A}}$. Specifically, the optimal occupancy measure corresponding to $\hat{r}$ should remain approximately optimal for $r^{\textsf{E}}$ under every transition law in $\mathcal{P}$. This yields the following definition of $\varepsilon$-transferability. {restatable}[$\varepsilon$-transferability]definitionepstransferability Fix some $\varepsilon>0$. We say that $\hat{r}$ is $\varepsilon$-transferable to some set of transition laws $\mathcal{P}\subseteq\Delta_{\mathcal{S}}^{\mathcal{S}\times\mathcal{A}}$ if $\ell_{P}(r^{\textsf{E}},\mathsf{RL}_{P}(\hat{r}))\leq\varepsilon$ for all $P\in\mathcal{P}$. We say that $\hat{r}$ is exactly transferable to $\mathcal{P}$ if it is $\varepsilon$-transferable to $\mathcal{P}$ with $\varepsilon=0$. The error margin of $\varepsilon$ is crucial, as exact transferability is unrealistic when learning from finite expert data. Moreover, note that Definition 5 is a definition of uniform transferability, as it requires $\hat{r}$ to be $\varepsilon$-transferable to any $P\in\mathcal{P}$ with the same fixed $\varepsilon$. In the following, we will analyze the transferability of a reward $\hat{r}$ for which all experts are $\hat{\varepsilon}$-optimal for some $\hat{\varepsilon}>0$. That is,
$\ell_{P^{k}}(\hat{r},\mu^{\textsf{E}}_{P^{k}})\leq\hat{\varepsilon},\quad k=0,\ldots,K-1.$ (6)
@@ 3.2 Related work
is satisfied (Rolland et al., 2022). Moreover, Schlaginhaufen and Kamgarpour (2023) show that identifying the expert’s reward up to a constant is a necessary and sufficient condition for exact transferability to any set $\mathcal{P}\subseteq\Delta_{\mathcal{S}}^{\mathcal{S}\times\mathcal{A}}$ whose interior, $\operatorname{\operatorname{int}}\mathcal{P}$, is non-empty and open (with respect to the subspace topology on $\Delta_{\mathcal{S}}^{\mathcal{S}\times\mathcal{A}}$ (Bourbaki, 1966)).
@@
We consider a two-state, two-action MDP with $\mathcal{S}=\mathcal{A}=\left\{{0,1}\right\}$, uniform initial state distribution, discount rate $\gamma=0.9$, and Shannon entropy regularization $h=-\mathcal{H}$ (see Appendix C). Suppose the expert reward is $r^{\textsf{E}}(s,a)=\mathbbm{1}\{s=1\}$ and consider the transition laws, $P^{0}$ and $P^{1}$, defined by $P^{0}(0|s,a)=1$ and $P^{1}(0|s,a)=\beta\cdot\mathbbm{1}\left\{{s=0,a=0}\right\}$ for some $\beta\in(0,1)$. Also, consider the two experts $\mu^{\textsf{E}}_{P^{0}}=\mathsf{RL}_{P^{0}}(r^{\textsf{E}})$ and $\mu^{\textsf{E}}_{P^{1}}=\mathsf{RL}_{P^{1}}(r^{\textsf{E}})$, and suppose we recovered the reward $\hat{r}(s,a)=-r^{\textsf{E}}$. Then, as detailed in Appendix E, the following holds: 1) We have $\ell_{P^{0}}(\hat{r},\mu^{\textsf{E}}_{P^{0}})=0$ and $\ell_{P^{1}}(\hat{r},\mu^{\textsf{E}}_{P^{1}})=\mathcal{O}(\beta)$. That is, for small $\beta$, the reward $\hat{r}$ is a good solution to the IRL problem, as both experts are approximately optimal under $\hat{r}$. 2) The rank condition (9) between $P^{0}$ and $P^{1}$ is satisfied for any $\beta\in(0,1)$. 3) For a new transition law $P$ defined by $P(0|s,a)=\mathbbm{1}\left\{{s=1,a=0}\right\}$, we have $\ell_{P}(r^{\textsf{E}},\mathsf{RL}_{P}(\hat{r}))\approx 4.81$, i.e. $\mathsf{RL}_{P}(\hat{r})$ performs poorly under the experts’ reward.

### v2

@@ 1 Introduction
  • We experimentally validate our results in a gridworld environment (Section 5). [^1] The code is openly accessible at https://github.com/andrschl/transfer_irl.
@@ 2 Background
Given $x$ and $y$ in some Euclidean vector space $\mathcal{V}$, we denote the $p$-norm by $\|x\|_{p}$, the orthogonal projection onto a closed convex set $\mathcal{X}\subset\mathcal{V}$ by $\Pi_{\mathcal{X}}(x)=\arg\min_{y\in\mathcal{X}}\|x-y\|_{2}$, and the standard dot product by $\langle x,y\rangle$. For a linear operator $A$, we denote its image and rank by $\operatorname{im}A$ and $\operatorname{\operatorname{rank}}A$, respectively. Given two sets $\mathcal{X}$ and $\mathcal{Y}$, we denote $\mathcal{X}+\mathcal{Y}$ for their Minkowski sum and $\mathcal{Y}^{\mathcal{X}}$ for the set of all functions mapping from $\mathcal{X}$ to $\mathcal{Y}$. Additionally, we denote $\Delta_{\mathcal{X}}$ for the probability simplex over $\mathcal{X}$ and $\mathbbm{1}$ for the indicator function. The interior $\operatorname{\operatorname{int}}\mathcal{X}$, the relative interior $\operatorname{\operatorname{relint}}\mathcal{X}$, the relative boundary $\operatorname{\operatorname{relbd}}\mathcal{X}$, and the convex hull $\operatorname{\operatorname{conv}}\mathcal{X}$ of some set $\mathcal{X}$ are defined in Appendix A, along with an overview of all other notations.
@@
We consider a regularized MDP (Geist et al., 2019) defined by a tuple $\left({\mathcal{S},\mathcal{A},P,\nu_{0},r,\gamma,h}\right)$. Here, $\mathcal{S}$ and $\mathcal{A}$ represent finite state and action spaces with $|\mathcal{S}|,|\mathcal{A}|>1$, $\nu_{0}\in\Delta_{\mathcal{S}}$ the initial state distribution, $P\in\Delta_{\mathcal{S}}^{\mathcal{S}\times\mathcal{A}}$ the transition law, $r\in\mathbb{R}^{\mathcal{S}\times\mathcal{A}}$ the reward, and $\gamma\in(0,1)$ the discount factor. Furthermore, $h:\mathcal{X}\to\mathbb{R}$ is a strictly convex regularizer that is defined on a closed convex set $\mathcal{X}\subseteq\mathbb{R}^{\mathcal{A}}$ with $\operatorname{\operatorname{relint}}\Delta_{\mathcal{A}}\subseteq\operatorname{\operatorname{int}}\mathcal{X}$. Starting from some initial state $s_{0}\sim\nu_{0}$ the agent can at each step in time $t$, choose an action $a_{t}\in\mathcal{A}$, will arrive in state $s_{t+1}\sim P(\cdot|s_{t},a_{t})$, and receives reward $r(s_{t},a_{t})$. The goal is to find a Markov policy $\pi\in\Delta_{\mathcal{A}}^{\mathcal{S}}$ maximizing the regularized objective $\mathbb{E}_{\pi}\left[{\sum_{t=0}^{\infty}\gamma^{t}\left[{r(s_{t},a_{t})-h\left({\pi(\cdot|s_{t})}\right)}\right]}\right]$. Following the classical linear programming approach to MDPs (Puterman, 2014), this can be cast equivalently as the convex optimization problem
@@
Due to the strict convexity of $h$, the regularized MDP problem has a unique optimal policy (Geist et al., 2019), hence guaranteeing the uniqueness of the optimal occupancy measure in (O-RL). In addition, we assume that the gradients of $h$ become unbounded towards the relative boundary of the simplex as detailed in Assumption 2 below. {restatable}[Steep regularization]assumptionsteepregularization Suppose that $h:\mathcal{X}\to\mathbb{R}$ is differentiable in $\operatorname{\operatorname{int}}\mathcal{X}$ and that $\lim_{l\to\infty}\left\lVert\nabla h(p_{l})\right\rVert_{2}=\infty$ if $\left({p_{l}}\right)_{l\in\mathbb{N}}$ is a sequence in $\operatorname{\operatorname{int}}\mathcal{X}$ converging to a point $p\in\operatorname{\operatorname{relbd}}\Delta_{\mathcal{A}}$. Assumption 2 ensures that the optimal policy is non-vanishing, and together with Assumption 2 below, we also have that the optimal occupancy measure is non-vanishing.
{restatable}
[Exploration]assumptionocclowerbound Let $\nu(s)\vcentcolon=\sum_{a}\mu(s,a)\geq\nu_{\text{min}}>0$ for any $s\in\mathcal{S}$ and $\mu\in\mathcal{M}$. One way to guarantee Assumption 2 is to impose a lower bound on the initial state distribution $\nu_{0}$. In the following, it will be convenient to denote the optimal solution to (O-RL) for the reward $r$ as
@@
$\mathsf{SubOpt}(r,\mu)\vcentcolon=\max_{\mu^{\prime}\in\mathcal{M}}J(r,\mu^{\prime})-J(r,\mu).$ (3)
That is, $\mu=\mathsf{RL}(r)$ if and only if $\mathsf{SubOpt}(r,\mu)=0$.
@@
As we aim to analyze the transferability of rewards to new transition laws $P\in\Delta_{\mathcal{S}}^{\mathcal{S}\times\mathcal{A}}$, it will often be useful to explicitly specify the dependency on $P$. We do so by adding a subscript – e.g. we write $\mathcal{M}_{P}$, $\mathsf{RL}_{P}$, and $\mathsf{SubOpt}_{P}$. However, for better readability, we drop these subscripts whenever there is no potential for confusion.
@@
However, since we don’t have direct access to the expert’s policy but only to a finite set of demonstrations, the best we can hope for is an algorithm that with high probability outputs a reward $\hat{r}\in\mathcal{R}$ such that $\mathsf{SubOpt}(\hat{r},\mu^{\textsf{E}})$ is small – i.e. an algorithm that is PAC (Syed and Schapire, 2007).
@@
The reward corresponding to an optimal occupancy is not unique. For example, all rewards in the affine subspace $r+\mathcal{U}$, where $\mathcal{U}\vcentcolon=\operatorname{im}(E-\gamma P)$ is the subspace of so-called potential shaping transformations, correspond to the same optimal occupancy measure (Ng et al., 1999). From a geometric perspective, the subspace $\mathcal{U}=\text{im}(E-\gamma P)$ lies perpendicular to the set of occupancy measures $\mathcal{M}$. Therefore, adding an element of $\mathcal{U}$ to the reward leaves the performance difference between any two occupancy measures invariant. Hence, it is often convenient to consider these rewards as equivalent (Kim et al., 2021) and to measure distances between rewards in the resulting quotient space. Given a linear subspace $\mathcal{V}\subset\mathbb{R}^{\mathcal{S}\times\mathcal{A}}$, the quotient space $\mathbb{R}^{\mathcal{S}\times\mathcal{A}}/\mathcal{V}$ is the set of all equivalence classes $[r]_{\mathcal{V}}\vcentcolon=\left\{{r^{\prime}\in\mathbb{R}^{\mathcal{S}\times\mathcal{A}}:r^{\prime}-r\in\mathcal{V}}\right\}$, which is itself a vector space with addition and multiplication operation defined by $[r]_{\mathcal{V}}+[r^{\prime}]_{\mathcal{V}}=[r+r^{\prime}]_{\mathcal{V}}$ and $c[r]_{\mathcal{V}}=[cr]_{\mathcal{V}}$ for $c\in\mathbb{R}$. Intuitively, $\mathbb{R}^{\mathcal{S}\times\mathcal{A}}/\mathcal{V}$ is the vector space obtained by collapsing $\mathcal{V}$ to zero, or in other words, it is isomorphic to the orthogonal complement of $\mathcal{V}$. We endow $\mathbb{R}^{\mathcal{S}\times\mathcal{A}}/\mathcal{V}$ with the quotient norm $\left\lVert[r]_{\mathcal{V}}\right\rVert_{2}\vcentcolon=\min_{v\in\mathcal{V}}\left\lVert r+v\right\rVert_{2}=\left\lVert\Pi_{\mathcal{V}^{\perp}}r\right\rVert_{2}$ and we say that $r$ and $r^{\prime}$ are close in $\mathbb{R}^{\mathcal{S}\times\mathcal{A}}/\mathcal{V}$ if $\left\lVert[r]_{\mathcal{V}}-[r^{\prime}]_{\mathcal{V}}\right\rVert_{2}$ is small. Moreover, the expert’s reward is said to be identifiable up to some equivalence class $[\cdot]_{\mathcal{V}}$ if $\mathsf{IRL}(\mu^{\textsf{E}})\subseteq[r^{\textsf{E}}]_{\mathcal{V}}$. In this paper, we will consider the equivalence relations induced by constant shifts, i.e., $\mathcal{V}=\bm{1}\vcentcolon=\left\{{r\in\mathbb{R}^{\mathcal{S}\times\mathcal{A}}:r(s,a)=c\in\mathbb{R}}\right\}$, and by potential shaping transformations, i.e., $\mathcal{V}=\mathcal{U}$. Note that since $\bm{1}$ is a subspace of $\mathcal{U}$ and $\mathcal{U}$ is $|\mathcal{S}|$-dimensional, $[r]_{\bm{1}}$ is a strict subset of $[r]_{\mathcal{U}}$ whenever $|\mathcal{S}|>1$.
@@ 3.1 Problem formulation
consisting of trajectories sampled independently from the experts $\mu^{\textsf{E}}_{P^{0}},\ldots,\mu^{\textsf{E}}_{P^{K-1}}$. Each expert is optimal for the same unrevealed reward $r^{\textsf{E}}\in\mathcal{R}$, but under different transition laws, $P^{0},\ldots,P^{K-1}$. Our goal is to recover a reward $\hat{r}\in\mathcal{R}$ that is transferable across a set of transition laws $\mathcal{P}\subseteq\Delta_{\mathcal{S}}^{\mathcal{S}\times\mathcal{A}}$. Specifically, the optimal occupancy measure corresponding to $\hat{r}$ should remain approximately optimal for $r^{\textsf{E}}$ under every transition law in $\mathcal{P}$. This yields the following definition of $\varepsilon$-transferability. {restatable}[$\varepsilon$-transferability]definitionepstransferability Fix some $\varepsilon>0$. We say that $\hat{r}$ is $\varepsilon$-transferable to some set of transition laws $\mathcal{P}\subseteq\Delta_{\mathcal{S}}^{\mathcal{S}\times\mathcal{A}}$ if $\mathsf{SubOpt}_{P}(r^{\textsf{E}},\mathsf{RL}_{P}(\hat{r}))\leq\varepsilon$ for all $P\in\mathcal{P}$. We say that $\hat{r}$ is exactly transferable to $\mathcal{P}$ if it is $\varepsilon$-transferable to $\mathcal{P}$ with $\varepsilon=0$. The error margin of $\varepsilon$ is crucial, as exact transferability is unrealistic when learning from finite expert data. Moreover, note that Definition 5 is a definition of uniform transferability, as it requires $\hat{r}$ to be $\varepsilon$-transferable to any $P\in\mathcal{P}$ with the same fixed $\varepsilon$. In the following, we will analyze the transferability of a reward $\hat{r}$ for which all experts are $\hat{\varepsilon}$-optimal for some $\hat{\varepsilon}>0$. That is,
$\mathsf{SubOpt}_{P^{k}}(\hat{r},\mu^{\textsf{E}}_{P^{k}})\leq\hat{\varepsilon},\quad k=0,\ldots,K-1.$ (6)
@@
Remark 3.1.
As discussed in Appendix J, the assumption of perfect expert optimality with respect to $r^{\textsf{E}}$ can be relaxed to allow for a misspecification error. All our results remain applicable in this setting but include an additional error term due to the experts’ suboptimality.
@@ 3.2 Related work
is satisfied (Rolland et al., 2022). Moreover, Schlaginhaufen and Kamgarpour (2023) show that identifying the expert’s reward up to a constant is a necessary and sufficient condition for exact transferability to any full-dimensional set $\mathcal{P}\subseteq\Delta_{\mathcal{S}}^{\mathcal{S}\times\mathcal{A}}$ (a set $\mathcal{P}$ whose interior, with respect to the subspace topology on $\Delta_{\mathcal{S}}^{\mathcal{S}\times\mathcal{A}}$ (Bourbaki, 1966), is non-empty).
@@
We consider a two-state, two-action MDP with $\mathcal{S}=\mathcal{A}=\left\{{0,1}\right\}$, uniform initial state distribution, discount rate $\gamma=0.9$, and Shannon entropy regularization $h=-\mathcal{H}$ (see Appendix C). Suppose the expert reward is $r^{\textsf{E}}(s,a)=\mathbbm{1}\{s=1\}$ and consider the transition laws, $P^{0}$ and $P^{1}$, defined by $P^{0}(0|s,a)=0.75$ and $P^{1}(0|s,a)=0.25+\beta\cdot\mathbbm{1}\left\{{s=0,a=0}\right\}$ for some $\beta\in[0,0.75]$. Also, consider the two experts $\mu^{\textsf{E}}_{P^{0}}=\mathsf{RL}_{P^{0}}(r^{\textsf{E}})$ and $\mu^{\textsf{E}}_{P^{1}}=\mathsf{RL}_{P^{1}}(r^{\textsf{E}})$, and suppose we recovered the reward $\hat{r}(s,a)=-r^{\textsf{E}}$. Then, as detailed in Appendix E, the following holds: 1) We have $\mathsf{SubOpt}_{P^{0}}(\hat{r},\mu^{\textsf{E}}_{P^{0}})=0$ and $\mathsf{SubOpt}_{P^{1}}(\hat{r},\mu^{\textsf{E}}_{P^{1}})=\mathcal{O}(\beta)$. That is, for small $\beta$, the reward $\hat{r}$ is a good solution to the IRL problem, as both experts are approximately optimal under $\hat{r}$. 2) The rank condition (9) between $P^{0}$ and $P^{1}$ is satisfied for any $\beta>0$. 3) For a new transition law $P$ defined by $P(0|s,a)=\mathbbm{1}\left\{{s=1,a=0}\right\}$, we have $\mathsf{SubOpt}_{P}(r^{\textsf{E}},\mathsf{RL}_{P}(\hat{r}))\approx 4.81$, i.e. $\mathsf{RL}_{P}(\hat{r})$ performs poorly under the experts’ reward.